payload:

Define one relation that depicts the following scenario: Find set of numbers, x and y, such that x raised to y is n, and y raised to x is the reciprocal of n.

payload:

I get that \(x^{y}=n\) and \(y^{x}=\frac{1}{n}\) are the two relations but I need to make it one that shows the possible values of x and y??

justjm:

It looks very difficult to do but once you realize the trick then it's easy. Use substitution. You know that \(x^{y}=n\) and \(y^{x}=\frac{1}{n}\), but you also know that the n is the same. Use the mathematical property of substitution. \(x^{y}=\frac{1}{y^{x}}\) Now it's just simplifying. \(x^{y}=y^{-x}\) \(\ln\left(x^{y}\right)=\ln\left(y^{-x}\right)\) \(y\ln x=-x\ln y\) \(\frac{y}{\ln y}=-\frac{x}{\ln x}\) and poof, there's your relation!!

justjm:

Probably they'll except your answer as this: \(y=-\frac{x\ln y}{\ln x}\)

payload:

Thanks so much man!!!!! I actually get it! Can I post more in the new questions?